# Frequency Decomposition-Driven Unsupervised Domain Adaptation for Remote Sensing Image Semantic Segmentation

### v1

@@ 5.2 Evaluation Metrics
where $TP$, $FP$, $TN$ and $FN$ are true positives, false positives, true negatives and false negatives, respectively. Furthermore, the mean F1 score (mF1) and the mean IoU (mIoU) for the main five classes apart from Clutter/Background are calculated to measure the average performance of the UDA methods following the common practice reported in the literature (Yan et al., 2019; Ma et al., 2023b; Yan et al., 2021).
In addition, the computational complexity of the proposed GLGAN and FD-GLGAN is evaluated using the following metrics. First, the floating point operation count (FLOPs) is employed to evaluate the model complexity. Second, the number of model parameters (M) and the memory footprint (MB) are utilized to evaluate the memory requirement. Finally, the frames per second (FPS) is used to evaluate the execution speed. For a computationally efficient method, its first three metrics are usually small, while its FPS value should be large.
For comparison purposes, we benchmark the proposed model against several state-of-the-art models trained for $50$ epochs, each containing $1000$ batches. All models are evaluated every $100$ batch before their performances are recorded during the training stage. Finally, each method’s best model corresponding to the highest segmentation performance is recorded.
Table 4: Task 1: Experimental results under the variations in geographic locations only: Quantitative comparison (%) with the state-of-the-art on the adaptation P-IRRG to V-IRRG. We show F1 for each category and the best values are bolded except Supervised method.
@@
Non-DA (Source-Only)
ResUNet++
76.93
70.79
40.19
7.06
56.70
63.03
50.33
37.13
DANN
83.07
75.93
54.15
55.38
69.73
71.82
67.66
52.24
AdasegNet
ResNet101
87.87
79.44
56.73
57.59
76.10
76.29
71.55
57.14
Advent
ResNet101
90.85
78.78
58.31
55.07
79.53
77.63
72.51
58.68
GANAI
ResNet101
87.13
77.20
52.48
58.78
74.38
73.52
69.99
55.29
TriADA
ResNet101
89.34
79.06
47.50
65.72
81.11
77.56
72.55
58.88
CCAGAN
ResNet101
89.02
78.10
59.29
65.42
79.10
76.90
74.19
60.09
MBATA-GAN
ResNet101
90.98
79.51
63.41
66.38
82.80
80.75
76.77
63.50
GLGAN
Swin-Base
91.95
80.31
67.71
65.89
84.18
81.97
78.01
65.04
Swin-Large
92.33
81.60
69.11
67.34
84.58
82.74
78.99
66.30
FD-GLGAN
Swin-Base
93.94
80.95
70.06 70.40 86.13 83.66 80.30 68.09
Non-DA (Supervised)
ResUNet++
94.84
90.15
78.25
80.16
90.47
89.66
86.77
77.20
5.3 Experimental settings
In our work, UNetformer (Wang et al., 2022b) was adopted as the generator, which integrated the Swin Transformer (Liu et al., 2021) pretrained on ImageNet (Deng et al., 2009) as its encoder. In our experiments, two versions of Swin Transformers, namely Swin-Base and Swin-Large, were applied for image feature extraction. Both Swin-Base and Swin-Large have four stages with the stack of $\left\{2,2,18,2\right\}$ SwinTBs with the embedding dimensions of Swin-Base and Swin-Large being $\left\{128,256,512,1024\right\}$ and $\left\{196,384,768,1536\right\}$, respectively. Furthermore, the number of the heads in Swin-Base and Swin-Large are $\left\{4,8,16,32\right\}$ and $\left\{6,12,24,48\right\}$, respectively. In addition, the window size for W-MHSA and SW-MHSA is $16\times 16$. For both decoder and GLDis, their decoder dimension is $256$, whereas the window size of the global-local attention is $8\times 8$.
All models in the experiments were implemented with the PyTorch framework on a single NVIDIA GeForce RTX 3090 GPU with $24$-GB RAM. To efficiently read the large RS patches, the sliding window method was adopted to collect the training samples dynamically. During the training stage, the size of the sliding window was set to $256\times 256$ with no overlapping, whereas the batch size was $10$. Furthermore, the generator was optimized with the Stochastic Gradient Descent (SGD) optimizer and the Nesterov acceleration with a momentum of $0.9$ and a weight decay of $5\times 10^{-4}$. The initial learning rate was set to $2.5\times 10^{-4}$, decreasing with a polynomial decay of power $0.9$. For GLDis, the Adam optimizer was deployed with the learning rate of $10^{-4}$ with a polynomial decay of power $0.9$. The momentum of Adam is set to $\left[0.9,0.99\right]$. Finally, the multi-level adversarial learning scheme was also adopted with parameters set to their default values derived from AdasegNet (Tsai et al., 2018) in which $\lambda^{seg}$ and $\lambda^{adv}$ were set to $1.0$ and $0.001$, respectively. $\lambda^{f}$ was set to $0.01$ according to the parameter analysis that will be presented in Sec. 5.7.
Figure 9: Task 1: qualitative performance comparison of the adaptation P-IRRG to V-IRRG with the size of $2048\times 2048$. (a) RGB images, (b) Ground Truth, (c) Source-only, (d) AdasegNet, (e) Advent, (f) GANAI, (g) TriADA, (h) CCAGAN, (i) MBATA-GAN, (j) GLGAN, (k) FD-GLGAN and (l) Supervised method.
5.4 Benchmarking
To confirm the effectiveness of the proposed methods, we performed comparison experiments on two non-DA methods, namely a deep model without domain adaptation (labeled as “Source-only”) and a supervised deep model (labeled as “Supervised”) trained with labeled data from the target domain. These two non-DA methods stand for the two extreme cases, with the “Supervised” model representing the optimal performance that any UDA method targets and the “Source-only” model standing for the performance lower bound. Furthermore, the classical ResUNet++ (Jha et al., 2019) is selected for these two non-DA methods by taking into account the balance between segmentation performance and computational complexity. In addition to the two non-UDA methods, some representative existing UDA methods were also selected for quantitative comparison. Table 3 summarizes the nine methods evaluated in our following experiments.
For the aforementioned methods, the same network architectures, training parameters, and evaluation methods were applied to make a fair comparison.
5.5 Results and Discussions
In this section, extensive experimental results are reported for the three domain adaptation tasks listed in Table 2.
Table 5: Task 2: Experimental results under the variations in both geographic locations and imaging mode: Quantitative comparison (%) with the state-of-the-art on the adaptation P-RGB to V-IRRG. We show F1 for each category and the best values are bolded except Supervised method.
@@ 5.5 Results and Discussions
Non-DA (Source-Only)
ResUNet++
64.48
0.04
5.26
1.57
58.21
35.71
25.91
18.43
DANN
73.20
62.46
41.50
54.94
64.39
61.85
59.30
42.94
AdasegNet
ResNet101
74.04
75.16
40.85
57.02
62.12
64.75
61.84
45.92
Advent
ResNet101
83.29
77.85
42.37
49.92
65.73
67.26
63.83
48.84
GANAI
ResNet101
76.87
56.25
41.50
57.93
66.07
61.81
59.72
43.57
TriADA
ResNet101
77.86
73.85
36.57
58.70
66.48
66.09
62.69
47.20
CCAGAN
ResNet101
85.60
75.27
36.31
63.24
66.03
68.38
65.25
50.58
MBATA-GAN
ResNet101
87.67
75.30
39.06
64.95
67.33
69.29
66.86
52.31
GLGAN
Swin-Base
87.71
68.37
58.59
66.68
73.73
73.21
71.02
55.98
Swin-Large
89.70
77.53
58.93
63.21
78.48
77.23
73.57
59.44
FD-GLGAN
Swin-Base
92.42
77.53
62.88 69.38 80.42 79.16 76.53 63.09
Non-DA (Supervised)
ResUNet++
94.84
90.15
78.25
80.16
90.47
89.66
86.77
77.20
Figure 10: Task 2: qualitative results of the adaptation P-RGB to V-IRRG with the size of $2048\times 2048$. (a) RGB images, (b) Ground Truth, (c) Source-only, (d) AdasegNet, (e) Advent, (f) GANAI, (g) TriADA, (h) CCAGAN, (i) MBATA-GAN, (j) GLGAN, (k) FD-GLGAN and (l) Supervised method.
Task 1 (P-IRRG to V-IRRG): The results for Task $1$ are summarized in Table 4. Inspection of Table 4 suggests that the “Source-Only” method showed the worst performance among all methods under consideration in terms of OA, mF1 and mIoU. Furthermore, all UDA methods demonstrated performance improvement of different degrees. For instance, the simplest DANN achieved improvements of $8.79\%$ in OA, $17.33\%$ in mF1 and $15.11\%$ in mIoU, which confirmed the necessity and effectiveness of the UDA approach. In particular, the baseline Advent showed improved OA, mF1, and mIoU of $77.63\%$, $72.51\%$ and $58.68\%$, respectively. Compared with the baseline, the proposed GLGAN achieved great improvements on all matrices. Specifically, the OA, mF1 and mIoU attained by GLGAN (Swin-Base) were $81.97\%$, $78.01\%$ and $65.04\%$, respectively, which amounts to an improvement of $4.34\%$, $5.50\%$ and $6.36\%$ as compared to Advent. In addition, GLGAN (Swin-Large) attained higher scores in OA, mF1 and mIoU of $82.74\%$, $78.99\%$ and $66.30\%$, respectively, by employing a more complex encoder. Among all the unsupervised methods examined, the proposed FD-GLGAN achieved the highest scores on all matrices with the smallest performance gap as compared to the “Supervised” non-DA method. It is easy to observe in Table 4 that the proposed FD-GLGAN attained the highest F1 score in four main categories, including $93.94\%$ for Buildings, $70.06\%$ for Low Vegetation, $70.40\%$ for Cars and $86.13\%$ for Impervious Surface, which amounts to improvement of $3.09\%$ for Buildings, $11.75\%$ for Low Vegetation, $15.33\%$ for Cars and $6.6\%$ for Impervious Surface. Upon comparing the results of GLGAN (Swin-Base) and FD-GLGAN (Swin-Base), it becomes evident that improvements are observed across all categories, suggesting that the HLFD module facilitates the learning of transferable features.
Finally, Fig. 9 presents the semantic segmentation results achieved by all methods under consideration. Visual inspection of Fig. 9 reveals that the proposed GLGAN and FD-GLGAN are more accurate in differentiating visually similar categories such as Trees and Low Vegetation. Furthermore, the segmentation boundary of ground objects provided by our method is much smoother, with fewer random prediction points than other UDA methods.
Task 2 (P-RGB to V-IRRG): Table 5 shows the results for Task $2$ that encounters a more severe domain shift problem than Task $1$. In addition to the discrepancies in the spatial resolution, the geographical location and other environmental factors, Task $2$ also has to learn image representations from the RGB bands before transferring the learned features to the IRRG bands. Notably, the “Source-only” method showed the worst performance, which is evidenced by its poor OA, mF1, and mIoU of $35.71\%$, $25.91\%$ and $18.43\%$, respectively. In sharp contrast, it is observed from Table 5 that the proposed GLGAN (Swin-Large) achieved OA of $77.23\%$, mF1 of $73.57\%$, and mIoU of $59.44\%$, which stands for an increase of $9.97\%$, $9.74\%$, and $10.60\%$ as compared to the corresponding performance of Advent, respectively. Further, the proposed FD-GLGAN (Swin-Base) achieved a higher performance with an OA of $79.16\%$, mF1 of $76.53\%$, and mIoU of $63.09\%$ than the GLGAN (Swin-Large). It proved the broad applicability of the proposed HLFD, which can steadily improve the performance of the original UDA method in different scenarios. Moreover, these results demonstrate that the proposed FD-GLGAN can effectively handle large domain shifts in Task $2$. These results demonstrate that the proposed GLGAN and FD-GLGAN can effectively handle large domain shifts in Task $2$.
Table 6: Task 3: Experimental results on domain shift in both geographic location and imaging mode: Quantitative comparison (%) with the state-of-the-art on the adaptation V-IRRG to P-RGB. We show F1 for each category and the best values are bolded except Supervised method.

### v2

@@ IV-B Evaluation Metrics
where $TP$, $FP$, $TN$ and $FN$ are true positives, false positives, true negatives and false negatives, respectively. Furthermore, the mean F1 score (mF1) and the mean IoU (mIoU) for the main five classes apart from Clutter/Background are calculated to measure the average performance of the UDA methods following the common practice reported in the literature [44, 21].
In addition, the computational complexity of the proposed GLGAN and De-GLGAN is evaluated using the following metrics. First, the giga floating point operation counts (GFLOPs) are employed to evaluate the model complexity. Second, the number of model parameters and the memory footprint are utilized to evaluate the memory requirement. Finally, the frames per second (FPS) is used to evaluate the execution speed. For a computationally efficient method, its first three metrics are usually small, while its FPS value should be large. Each method undergoes training for $50$ epochs, with each epoch consisting of $1000$ batches. During training, the performance of each method is evaluated every 100 batches and recorded. The best-performing model for each method is then selected for subsequent testing based on its highest segmentation performance.
IV-C Experimental settings
In our work, the basic unit of the encoding block in the generator is adopted from the Swin Transformer [40] pretrained on ImageNet. In the experiments, two versions of Swin Transformers, namely Swin-Base and Swin-Large, were applied for image feature extraction. Both Swin-Base and Swin-Large have four stages with the stack of $\left\{2,2,18,2\right\}$ SwinTBs with the embedding dimensions of Swin-Base and Swin-Large being $\left\{128,256,512,1024\right\}$ and $\left\{196,384,768,1536\right\}$, respectively. Furthermore, the number of the heads in Swin-Base and Swin-Large are $\left\{4,8,16,32\right\}$ and $\left\{6,12,24,48\right\}$, respectively. In addition, the window size for W-MHSA and SW-MHSA is $16\times 16$. The dimension of both the decoder and GLDis is set to $256$, whereas the window size of the global-local attention is $8\times 8$.
IV-D Benchmarking
To confirm the effectiveness of the proposed methods, some representative existing UDA methods were selected for quantitative comparison. Table III summarizes the nine methods evaluated in our following experiments. Notably, in addition to the first three representative methods in the field of computer vision, the other six methods are specifically designed for remote sensing images. Furthermore, we also performed comparison experiments on two non-DA methods for the first three UDA tasks on the ISPRS Potsdam and Vaihingen dataset, namely a deep model without domain adaptation (labeled as “Source-only”) and a supervised deep model (labeled as “Supervised”) trained with labeled data from the target domain. These two non-DA methods stand for the two extreme cases, with the “Supervised” model representing the optimal performance that any UDA method targets and the “Source-only” model standing for the performance lower bound. Furthermore, the classical ResUNet++ [65] is selected for these two non-DA methods by taking into account the balance between segmentation performance and computational complexity. For the aforementioned methods, the same network architectures, training parameters, and evaluation methods were applied to make a fair comparison.
TABLE IV: Task 1: Quantitative comparison of F1 (%) on the adaptation of P-IRRG to V-IRRG.
@@ IV-D Benchmarking
Non-DA (Source-Only) ResUNet++ 76.93 70.79 40.19 7.06 56.70 63.03 50.33 37.13
DANN - 83.07 75.93 54.15 55.38 69.73 71.82 67.66 52.24
AdasegNet ResNet101 87.87 79.44 56.73 57.59 76.10 76.29 71.55 57.14
Advent ResNet101 90.85 78.78 58.31 55.07 79.53 77.63 72.51 58.68
GANAI ResNet101 87.13 77.20 52.48 58.78 74.38 73.52 69.99 55.29
TriADA ResNet101 89.34 79.06 47.50 65.72 81.11 77.56 72.55 58.88
CCAGAN ResNet101 89.02 78.10 59.29 65.42 79.10 76.90 74.19 60.09
MBATA-GAN ResNet101 90.98 79.51 63.41 66.38 82.80 80.75 76.77 63.50
MASNet ResNet101 90.13 80.90 59.11 66.70 81.44 79.87 75.66 62.13
DSSFNet ResNet101 92.05 80.76 60.44 66.29 82.68 80.67 76.45 63.27
GLGAN Swin-Base 91.95 80.31 67.71 65.89 84.18 81.97 78.01 65.04
Swin-Large 92.33 81.60 69.11 67.34 84.58 82.74 78.99 66.30
De-GLGAN Swin-Base 93.94 80.95 70.06 70.40 86.13 83.66 80.30 68.09
Non-DA (Supervised) ResUNet++ 94.84 90.15 78.25 80.16 90.47 89.66 86.77 77.20
IV-E Results and Discussions
Task 1 (P-IRRG to V-IRRG): The results for Task $1$ are summarized in Table IV. Inspection of Table IV suggests that the “Source-Only” method showed the worst performance among all methods under consideration in terms of OA, mF1 and mIoU. Furthermore, all UDA methods demonstrated performance improvement of different degrees. For instance, the simplest DANN achieved improvements of $8.79\%$ in OA, $17.33\%$ in mF1 and $15.11\%$ in mIoU, which confirmed the necessity and effectiveness of the UDA approach. In particular, the baseline Advent showed improved OA, mF1, and mIoU of $77.63\%$, $72.51\%$ and $58.68\%$, respectively. Compared with the baseline, the proposed GLGAN achieved great improvements in all performance metrics. Specifically, the OA, mF1 and mIoU attained by GLGAN (Swin-Base) were $81.97\%$, $78.01\%$ and $65.04\%$, respectively, which amounts to an improvement of $4.34\%$, $5.50\%$ and $6.36\%$ as compared to Advent. In addition, GLGAN (Swin-Large) attained higher scores in OA, mF1 and mIoU, by employing a more complex encoder. Among all the unsupervised methods examined, the proposed De-GLGAN achieved the highest scores on all overall matrices with the smallest performance gap as compared to the “Supervised” non-DA method. It is easy to observe in Table IV that the proposed De-GLGAN attained the highest F1 score in four main categories, including $93.94\%$ for Building, $70.06\%$ for Low Vegetation, $70.40\%$ for Cars and $86.13\%$ for Impervious Surface, which amounts to improvement of $3.09\%$ for Building, $11.75\%$ for Low Vegetation, $15.33\%$ for Cars and $6.6\%$ for Impervious Surface. Upon comparing the results of GLGAN (Swin-Base) and De-GLGAN (Swin-Base), it becomes evident that improvements are observed across all categories, suggesting that the HLFD module facilitates the learning of transferable features. Finally, Fig. 10 presents the semantic segmentation results achieved by part of the comparative methods. Visual inspection of Fig. 10 reveals that the proposed GLGAN and De-GLGAN are more accurate in differentiating visually similar categories such as Trees and Low Vegetation. Furthermore, the segmentation boundary of ground objects provided by our method is much smoother, with fewer random prediction points than other UDA methods.
TABLE V: Task 2: Quantitative comparison of F1 (%) on the adaptation of P-RGB to V-IRRG.
@@ IV-E Results and Discussions
Non-DA (Source-Only) ResUNet++ 64.48 0.04 5.26 1.57 58.21 35.71 25.91 18.43
DANN - 73.20 62.46 41.50 54.94 64.39 61.85 59.30 42.94
AdasegNet ResNet101 74.04 75.16 40.85 57.02 62.12 64.75 61.84 45.92
Advent ResNet101 83.29 77.85 42.37 49.92 65.73 67.26 63.83 48.84
GANAI ResNet101 76.87 56.25 41.50 57.93 66.07 61.81 59.72 43.57
TriADA ResNet101 77.86 73.85 36.57 58.70 66.48 66.09 62.69 47.20
CCAGAN ResNet101 85.60 75.27 36.31 63.24 66.03 68.38 65.25 50.58
MBATA-GAN ResNet101 87.67 75.30 39.06 64.95 67.33 69.29 66.86 52.31
MASNet ResNet101 85.31 77.79 39.69 51.92 74.44 71.59 65.83 51.43
DSSFNet ResNet101 87.15 77.66 48.77 56.02 73.46 73.14 68.61 53.98
GLGAN Swin-Base 87.71 68.37 58.59 66.68 73.73 73.21 71.02 55.98
Swin-Large 89.70 77.53 58.93 63.21 78.48 77.23 73.57 59.44
De-GLGAN Swin-Base 92.42 77.53 62.88 69.38 80.42 79.16 76.53 63.09
Non-DA (Supervised) ResUNet++ 94.84 90.15 78.25 80.16 90.47 89.66 86.77 77.20
Figure 10: Task 1: qualitative performance comparison of the adaptation P-IRRG to V-IRRG with images of size $1024\times 1024$, where the two samples are distinguished by subscripts. (a) IRRG images, (b) Ground Truth, (c) AdasegNet, (d) Advent, (e) TriADA, (f) CCAGAN, (g) MBATA-GAN (h) MASNet, (i) DSSFNet and (j) De-GLGAN. Purple boxes are marked to highlight the differences.
Figure 11: Task 2: qualitative results of the adaptation P-RGB to V-IRRG with the size of $1024\times 1024$, where the two samples are distinguished by subscripts. (a) IRRG images, (b) Ground Truth, (c) AdasegNet, (d) Advent, (e) TriADA, (f) CCAGAN, (g) MBATA-GAN (h) MASNet, (i) DSSFNet and (j) De-GLGAN. Purple boxes are marked to highlight the differences.
Task 2 (P-RGB to V-IRRG): Table V shows the results for Task $2$ that encounters a more severe domain shift problem than Task $1$. In addition to the discrepancies in the spatial resolution, the geographical location and other environmental factors, Task $2$ also has to learn image representations from the RGB bands before transferring the learned features to the IRRG bands. Notably, the “Source-only” method showed the worst performance, which is evidenced by its poor OA, mF1, and mIoU of $35.71\%$, $25.91\%$ and $18.43\%$, respectively. In sharp contrast, it is observed from Table V that the GLGAN (Swin-Large) achieved OA of $77.23\%$, mF1 of $73.57\%$, and mIoU of $59.44\%$, which stands for an increase of $9.97\%$, $9.74\%$, and $10.60\%$ as compared to the corresponding performance of Advent, respectively. Further, the proposed De-GLGAN (Swin-Base) achieved a higher performance than the GLGAN (Swin-Large). It proved the broad applicability of the proposed HLFD, which can steadily improve the performance of the original UDA method in different scenarios. Moreover, these results demonstrate that the proposed GLGAN and De-GLGAN can effectively handle large domain shifts. Furthermore, it is interesting to observe that the De-GLGAN improved the accuracy for Low Vegetation by $20.51\%$, but suffered from performance degradation of $0.32\%$ for Tree. This observation could be explained by the fact that the trees in Potsdam are mostly sparse deciduous trees, while those in Vaihingen have denser foliage and canopy, as shown in Fig. 8. As a result, Tree and Low Vegetation in Potsdam share similar spectral characteristics with Low Vegetation in Vaihingen. Finally, Fig. 11 illustrates some segmentation examples obtained in Task $2$ by various methods. It is clearly observed that the “Source-only” method suffered from the worst performance. Furthermore, despite the fact that existing UDA methods showed improvements of different degrees, they incurred deficiencies in the aspects of providing clear boundaries and differentiating visually similar categories. By leveraging its powerful global-local feature extraction capability and frequency decomposition frequency technology, the proposed De-GLGAN achieved segmentation performance close to that of the supervised non-DA method.
TABLE VI: Task 3: Quantitative comparison of F1 (%) on the adaptation of V-IRRG to P-RGB.